OpenStudy (anonymous):

Simplify the number using the imaginary unit i. Square root - 28

OpenStudy (zehanz):

First, write \[\sqrt{-28}=\sqrt{-1} \sqrt{28}\]Now, can you simplify sqrt(28)?

OpenStudy (anonymous):

the problem is i have no calculator.

OpenStudy (zehanz):

You don't need one! You have to simplify, not approximate. \[\sqrt{28}=\sqrt{4*7}\]Does this ring a bell?

OpenStudy (anonymous):

that means 28 = 28

OpenStudy (anonymous):

so the answer is i squareroot 28

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (zehanz):

No, \[\sqrt{28}=\sqrt{4*7}=\sqrt{4}*\sqrt{7}=2\sqrt{7}\] Because you also have \[\sqrt{-1}=i\]The answer is \[2i \sqrt{7}\]

OpenStudy (zehanz):

Well Geckoh2o, I surely don't hope it is, otherwise you could just type it in without understanding what is going on!